OpenStudy (anonymous):

Helppppp! giving medals! A sandbag was thrown downward from a building. The function f(t) = -16t2 - 32t + 128 shows the height f(t), in feet, of the sandbag after t seconds. Part A: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. (4 points) Part B: Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? (4 points) Part C: Use your answer in part B to determine the axis of symmetry for f(x)? (2 points)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

c

OpenStudy (helder_edwin):

A.- \[ \large f(t)=-16t^2-32t+128=-16(t^2+2t-8)=-16(t+4)(t-2) \]

OpenStudy (helder_edwin):

B.- \[ \large y=f(x)=-16t^2-32t+128 \] \[ \large y-128-16=-16(t^2+2t+1) \] \[ \large y-144=-16(t+1)^2 \]